OpenStudy (anonymous):

PLEASE HELP ! For each of the following pairs of orbitals in a multi-electron atom, indicate which is higher in energy: (a) 1s, 2s; (b) 3s, 3d; (c) 4f, 5s.

OpenStudy (chmvijay):

n+l rule will help u to find this :)

OpenStudy (anonymous):

So would in part a) 2s be bigger than 1s ? b) 3d< 3s C) 4f < 5s

OpenStudy (chmvijay):

find out for each orbit the n+l value :) example 2S in this case n = 2 l=0 n+l =2+0=2 and for 1s in case n= 1 and l=0 n+l =1+0 so 2S>1S like this find for all and find ur answer